RC meeting and Run Plan, Friday Sept. 29, 2023
Return to the RC Meeting Calendar
- 1 RC Meeting Information
- 2 Useful Links
- 3 Procedures
- 4 Status Update
RC Meeting Information
Time: Monday - Friday at 3:30pm
- Location: Counting house conference room
- ZoomGov Meeting Connection Information
Hall A call-in instructions:
- Call RC first.
- Techs on site 7am-4:30pm Mon-Thu, 7-11am Fri. RC will call Hall A work coordinator (Zach) during normal working hours. After hours, RC calls Tech on-call.
- After 9pm everything non-emergency needs to be deferred. What is emergency and what's not? Call RC. RC talks with Mark to decide.
Hall Access procedure:
- Access check list (binder on the counting room desk) needs to be filled out before making an access.
- Scan (take pictures with phone) and post it to the logbook.
- Any work inside the SBS, BB boundary need to have PS locked out -- requires two QEWs to throw the big switch. Exception: field measurements.
New instruction for shift crew when changing target:
- Beam restoration instructions have been updated. Printed version is in the counting room.
- When we change target, make sure the SOFs of Target BPMs are set to the correct values.
- SOFs for different targets: Needs an update after restoring beam tonight (Note: SOFs for major target positions are on the Control Room whiteboard.)
After speaking to the Crew Chief and PD, I was told that the decision had been made not to deliver beam to Hall A last night. Given this, I called off shifts for the SL on SWING and OWL (the TO had to stay to monitor alarms). During OWL, Hunter took EPR measurements for target cell calibrations, and did a few other target studies whilst the beam was off. Beam should have been restored during today's DAY shift, but the Ion Chambers are being tripped whenever MCC attempts to deliver beam (https://logbooks.jlab.org/entry/4189307), so this hasn't been done yet. We are still hoping that beam will be back by the end of shift.
A Lenovo power supply was brought out of the hall ysterday (and has been cleared by RadCon). If this is yours, it's in the Counting House.
- Those responsible for looking at the PSS had been working on the assumption that the BSY PSS trips were something to do with the Hall A beamline coming out of the switchyard.
- They made the decision to not deliver beam to Hall A so that at least Hall C could take data without the risk of the PSS killing the accelerator overnight. This decision was not initially communicated to the PD, and therefore, I only found out when I inquired with the Crew Chief (who had been told).
- Overnight, the BSY PSS tripped anyway, indicating that the Hall A line was not responsible for the drops as first thought, so MCC agreed to deliver beam to Hall A again today.
- Beam recovery only started at noon today because of an issue with one of the South Linac segments which took 5 hours to fix.
- OPS still want to put a bypass in the PSS, so that halls can get beam again without the PSS trips dropping the accelerator; this is going through the proper channels, and will hopefully be installed by the end of today.
- No maintenance day on the MCC white board (until 10/10/2023).
Data collected :
- Accumulated corrected charge collected: 0C (no beam = no charge...) ...again.
Hall A Subsystems Status
- Latest target polarization: 37.1% [10:57 am]
- BBHodo: One of the channels (49R) is reporting an issue with the create connection. The set point voltage for this has been put to 0V.
- HCAL: Alex went into the hall to replace the fADC module that he and Jiwan looked into fixing on Wednesday.
Hall access requests
- If anyone else wants to make an access, let me know! There is no planned maintenance day before Tuesday 10th, so any accesses need to be opportunistic.
- Still need to fill open slots
- Open slots on Saturday 30th: DAY (TO).
Please sign up for shifts if there are any open!
Short Term Run plan
- Beam recovery to the hall
- Production run on polarized He3 target @35uA beam current.
- If MCC are restricting us to <35uA, then run at the maximum allowed, increasing it whenever MCC is happy.
- If the beam and target are stable at 35uA, call the RC. We can then go up to 40uA.
==> List of items in case He3 production running isn’t ready:
- Take data on reference cell at 10 atm H2 (portions that add up to 1 day) [85% (4200 elastic events) more data required for Bigbite momentum calibration]
- BigBite Sieve slit => only during day (Techs needed)
- with 8 carbon foils and take optics for 1 hr at 40 uA (Sean/Andrew - confirm time)
- 1 hour for single foil (magnets off) => Chance BBcal HV setting to appropriate values
- BigBite/single foil location reconstruction (~20 mins?) - BigBite + HCAL detector experts on call consulting analysis experts like Andrew/Sean (magnet on)